OpenStudy (anonymous):

Find an expression for the function whose graph is the given curve. The top half of the circle x^2 + (y − 3)^2 = 4 I am confused with what this question is asking.

OpenStudy (anonymous):

well, what is asking is to find a function that defines the top half of this circle

OpenStudy (anonymous):

Ok thanks, I understand what it is trying to say now.

OpenStudy (anonymous):

now to do so, we must first solve this expression for y. We can accomplish this by: (y-3)^2=4-x^2 Now, lets take the square root of both sides: y-3=±sqrt(4-x^2) y=±sqrt(4-x^2)+3 Now, we only need the top half or the positive half so this means that our answer is simply: y=sqrt(4-x^2)+3

OpenStudy (anonymous):

Thanks but next time if you decide to help me, please don't show a solution. I would just like hints.

OpenStudy (anonymous):

my bad, just cover your eyes :)